OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x))=x and g(f(x))=x. f(x)=(x-7)/(x+3) g(x)=(-3x-7)/(x-1)

OpenStudy (anonymous):

Complicated but doable. Sometimes can show A and A' are inverses by showing A*A' = 1. That does not fit the instructions, but could be interesting. (x-7)*(x-1)/(x+3)(-3x-7) = (x^2 -8x +7)/(-3x^2-16x-21) well, that certainly does not = 1 !

OpenStudy (anonymous):

lol so what am I supposed to do?

OpenStudy (anonymous):

Go through the tedious process of substituting f(x) into each x in g(x). The substitute g(x) for each x in f(x). and you should find g(f(x))=x and f(g(x))=x if the hypothesis and instructions are true.

OpenStudy (anonymous):

Everytime I try I end up with some crazy stuff. I start out like this: (x-7)/(x+3) and try to change replace the x's with the g(x) [(-3x-7-7)/(x-1)]/[(-3x-7+3)/(x-1) and it gets hectic from there

OpenStudy (anonymous):

It does get hectic. I'm glad it's someone else's problem! Good luck.

OpenStudy (anonymous):

Lol gee thanks.